OpenStudy (anonymous):

Solve C = 3(2f – 5h) for f in terms of C and h. You must show all work.

OpenStudy (anonymous):

shoot, thats wrong

OpenStudy (anonymous):

C=3(2f-5h) C/3=2f-5h C/3+5h=2f 1/2*(C/3+5h)=f